OpenStudy (anonymous):

find the LCD 5 , -3 = 18x^2 y^3 24x^4 y^5

OpenStudy (mertsj):

Write the prime factor of 18 and 24 in exponential form.

OpenStudy (anonymous):

18 = 3^2 * 2 24 = 2^3 * 3

OpenStudy (mertsj):

Now take the largest power of every DIFFERENT factor. That will be the LCM of the number part.

OpenStudy (mertsj):

And then do the same for the variable part and that will be your LCD

OpenStudy (anonymous):

2^3 * 3^2*x^4*y^5

OpenStudy (mertsj):

Good Job!!